A boy is pulled 20m along a horizontal plane with a constant force of 10 N, applied (I) parallel to the plane( II) in direction of angle 60° to the horizontal calculate the work done in each case


1
Expert's answer
2021-02-11T14:07:22-0500

"W_1=Fs\cos\alpha=10\cdot20\cdot\cos0°=200(J)" . Answer


"W_2=Fs\cos\alpha=10\cdot20\cdot\cos60°=100(J)" . Answer